# On the Efficacy of Text-Based Input Modalities for Action Anticipation

### v2

@@ 3.1 Pre-training
They are converted into a sentence using the template: A video containing the following
objects: <list of objects>
, and encoded using the aforementioned CLIP text encoder;
@@
We convert the class names into sentences using the prompt: Describe
<xyz> action in 1 sentence in 10
different ways
, and randomly select one response during training.
@@ 4.3 Ablations and Analysis
(ii) w/o gpt involves pre-training with the simple template - This is a
video clip with action
<xyz>
@@ Appendix B Contribution of Each Modality on Action Anticipation
1 | ‘sponge, tap’, ‘sponge, tap’,‘sponge, tap’,‘sponge, tap’, ‘sponge, tap’,‘sponge, tap’,‘sponge, tap’, ‘sponge, tap’,‘sponge, tap’,‘sponge, tap’ | ‘wash plate’, ‘wash plate’, ‘no action’, ‘wash plate’, ‘wash plate’, ‘wash plate’, ‘wash plate’, ‘wash plate’,‘insert plate’, ‘insert plate’, ‘wash sponge’,‘wash sponge’,‘wash sponge’, ‘wash sponge’,‘wash sponge’, ‘wash sponge’ | Wash cloth |
2 | ‘bin, spoon’, ‘bin’, ‘knife, ’, ‘knife, ’, ‘bin’, ‘bin’, ‘bag’, ‘bag’, ‘bag’, ‘bin, bag’, ‘bin, bag’,‘bin, bag’, ‘bin, bag’, ‘bin, bag’, ‘bin, bag’, ‘bin, bag’ | ‘wrap bag’, ‘wrap bag’, ‘wrap bag’, ‘wrap bag’,‘wrap bag’, ‘wrap bag’, ‘wrap bag’, ‘wrap bag’,‘wrap bag’, ‘wrap bag’, ‘wrap bag’, ‘wrap bag’,‘wrap bag’, ‘wrap bag’, ‘wrap bag’ | Tie Bag |
3 | ‘cupboard’, ‘cupboard’, ‘cupboard’, ‘cupboard’, ‘cupboard’, ‘pan,cupboard’,‘pan,cupboard’, ‘cupboard’, ‘cupboard’, ‘cupboard, lid’, ‘pan,cupboard’, ‘pan,cupboard’,‘pan,cupboard’, ‘pan,cupboard’, ‘pan,cupboard’, ‘pan, ’ | ‘take plate’, ‘take plate’, ‘take plate’, ‘take plate’, ‘take plate’, ‘no action’, ‘open cupboard’, ‘no action’, ‘insert plate’, ‘insert plate’, ‘no action’, ‘no action’, ‘take cup’,‘no action’, ‘open cupboard’, ‘insert cup’ | Put-into Cup |
@@
4 | ‘bowl,spoon, tap, knife’, ‘bowl,spoon, tap, knife’, ‘bowl,spoon, tap, knife’,‘bowl, spoon, cup, tap, knife’, ‘bowl, spoon, tap, knife’, ‘bowl, spoon, tap, knife’,‘bowl, spoon, cup, knife, bottle’, ‘bowl, cup, tap, knife, lid’, ‘bowl,knife, tap’,‘bowl, spoon, tap, knife, lid’, ‘bowl, spoon, tap, knife’, ‘bowl,knife, tap’,‘bowl, spoon, tap, knife’, ‘bowl, spoon, tap, knife, sponge’, ‘knife,tap, spoon’ | ‘wash cup’, ‘wash cup’, ‘no action’, ‘wash spoon’, ‘wash spoon’, ‘put spoon’, ‘wash cup’, ‘wash cup’, ‘wash cup’,‘wash cup’, ‘wash cup’, ‘wash cup’, ‘wash cup’, ‘no action’, ‘no action’, ’ turn-off tap’ | Turn-off tap |
@@
5 | ‘board:chopping, onion, knife, spatula’, ‘board:chopping, onion, knife, spatula’, ‘knife, onion, spatula’, ‘knife, board:chopping, onion, spatula’, ‘food, onion, knife, spatula’,‘knife, board:chopping, onion, spatula’, ‘knife, board:chopping, onion, spatula’,‘knife, board:chopping,spatula’, ‘knife, board:chopping, lid, spatula’, ‘board:chopping,knife’, ‘knife, board:chopping,spatula’, | ‘put board:chopping’, ‘no action’, ‘no action’, ‘put knife’, ‘no action’, ‘open drawer’, ‘no action’, ‘no action’,‘take spatula’, ‘close drawer’, ‘no action’, ‘mix aubergine’, ‘mix aubergine’, ‘mix aubergine’, ‘mix aubergine’, ‘put aubergine’ | Take salt |
@@
6 | ‘lid, glass, bottle’, ’lid’, ’lid, glass’, ’glass, lid’, ‘glass, lid’, ‘glass, container’, ‘glass, bottle, container’, ‘bag, glass’, ‘bag, glass’, ‘glass, bag, bottle’, ‘glass, bottle, oil’,’ glass, bag, bottle, container’, ‘glass, bag, lid, bottle’, spoon, bottle, glass, bag, lid’, ‘onion, lid’, ‘cup, glass, bottle, bag, lid’ | ‘move bin’, ’ move bin’, ‘take milk’, ‘take milk’,’ crush milk’, ‘crush milk’, ‘crush milk’, ‘crush milk’,‘no action’, ‘no action’, ‘no action’, ’ insert milk’,‘no action’, ‘no action’, ‘take paper’, ‘take paper’ | Move glass |

### v3

@@ 3.1 Pre-training
They are converted into a sentence using the template: A video containing the following objects: <list of objects>, and encoded using the aforementioned CLIP text encoder;
@@
We convert the class names into sentences using the prompt: Describe <xyz> action in 1 sentence in 10
different ways, and randomly select one response during training.
@@ 4.3 Ablations and Analysis
(ii) w/o gpt involves pre-training with the simple template - This is a video clip with action <xyz>;
@@ Appendix B Contribution of Each Modality on Action Anticipation
| 1 | ‘sponge, tap’, ‘sponge, tap’,‘sponge, tap’,‘sponge, tap’, ‘sponge, tap’,‘sponge, tap’,‘sponge, tap’, ‘sponge, tap’,‘sponge, tap’,‘sponge, tap’ | ‘wash plate’, ‘wash plate’, ‘no action’, ‘wash plate’, ‘wash plate’, ‘wash plate’, ‘wash plate’, ‘wash plate’,‘insert plate’, ‘insert plate’, ‘wash sponge’,‘wash sponge’,‘wash sponge’, ‘wash sponge’,‘wash sponge’, ‘wash sponge’ | Wash cloth |
| 2 | ‘bin, spoon’, ‘bin’, ‘knife, ’, ‘knife, ’, ‘bin’, ‘bin’, ‘bag’, ‘bag’, ‘bag’, ‘bin, bag’, ‘bin, bag’,‘bin, bag’, ‘bin, bag’, ‘bin, bag’, ‘bin, bag’, ‘bin, bag’ | ‘wrap bag’, ‘wrap bag’, ‘wrap bag’, ‘wrap bag’,‘wrap bag’, ‘wrap bag’, ‘wrap bag’, ‘wrap bag’,‘wrap bag’, ‘wrap bag’, ‘wrap bag’, ‘wrap bag’,‘wrap bag’, ‘wrap bag’, ‘wrap bag’ | Tie Bag |
| 3 | ‘cupboard’, ‘cupboard’, ‘cupboard’, ‘cupboard’, ‘cupboard’, ‘pan,cupboard’,‘pan,cupboard’, ‘cupboard’, ‘cupboard’, ‘cupboard, lid’, ‘pan,cupboard’, ‘pan,cupboard’,‘pan,cupboard’, ‘pan,cupboard’, ‘pan,cupboard’, ‘pan, ’ | ‘take plate’, ‘take plate’, ‘take plate’, ‘take plate’, ‘take plate’, ‘no action’, ‘open cupboard’, ‘no action’, ‘insert plate’, ‘insert plate’, ‘no action’, ‘no action’, ‘take cup’,‘no action’, ‘open cupboard’, ‘insert cup’ | Put-into Cup |
@@
| 4 | ‘bowl,spoon, tap, knife’, ‘bowl,spoon, tap, knife’, ‘bowl,spoon, tap, knife’,‘bowl, spoon, cup, tap, knife’, ‘bowl, spoon, tap, knife’, ‘bowl, spoon, tap, knife’,‘bowl, spoon, cup, knife, bottle’, ‘bowl, cup, tap, knife, lid’, ‘bowl,knife, tap’,‘bowl, spoon, tap, knife, lid’, ‘bowl, spoon, tap, knife’, ‘bowl,knife, tap’,‘bowl, spoon, tap, knife’, ‘bowl, spoon, tap, knife, sponge’, ‘knife,tap, spoon’ | ‘wash cup’, ‘wash cup’, ‘no action’, ‘wash spoon’, ‘wash spoon’, ‘put spoon’, ‘wash cup’, ‘wash cup’, ‘wash cup’,‘wash cup’, ‘wash cup’, ‘wash cup’, ‘wash cup’, ‘no action’, ‘no action’, ’ turn-off tap’ | Turn-off tap |
@@
| 5 | ‘board:chopping, onion, knife, spatula’, ‘board:chopping, onion, knife, spatula’, ‘knife, onion, spatula’, ‘knife, board:chopping, onion, spatula’, ‘food, onion, knife, spatula’,‘knife, board:chopping, onion, spatula’, ‘knife, board:chopping, onion, spatula’,‘knife, board:chopping,spatula’, ‘knife, board:chopping, lid, spatula’, ‘board:chopping,knife’, ‘knife, board:chopping,spatula’, | ‘put board:chopping’, ‘no action’, ‘no action’, ‘put knife’, ‘no action’, ‘open drawer’, ‘no action’, ‘no action’,‘take spatula’, ‘close drawer’, ‘no action’, ‘mix aubergine’, ‘mix aubergine’, ‘mix aubergine’, ‘mix aubergine’, ‘put aubergine’ | Take salt |
@@
| 6 | ‘lid, glass, bottle’, ’lid’, ’lid, glass’, ’glass, lid’, ‘glass, lid’, ‘glass, container’, ‘glass, bottle, container’, ‘bag, glass’, ‘bag, glass’, ‘glass, bag, bottle’, ‘glass, bottle, oil’,’ glass, bag, bottle, container’, ‘glass, bag, lid, bottle’, spoon, bottle, glass, bag, lid’, ‘onion, lid’, ‘cup, glass, bottle, bag, lid’ | ‘move bin’, ’ move bin’, ‘take milk’, ‘take milk’,’ crush milk’, ‘crush milk’, ‘crush milk’, ‘crush milk’,‘no action’, ‘no action’, ‘no action’, ’ insert milk’,‘no action’, ‘no action’, ‘take paper’, ‘take paper’ | Move glass |